OpenStudy (anonymous):

what is equivalent to log 25 125 ?

OpenStudy (anonymous):

125 log 25 maybe? or do you mean log base 25 of 125?

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

ok then, what exponent do you need to put over the 25 to get 125?

OpenStudy (anonymous):

25^0 = 1 25^1=25 25^2=625 so the answer is between 1 and 2